An electric vehicle starts from rest and accelerates at a rate of 2.2 m/s2 in a straight line until it reaches a speed of 22 m/s. The vehicle then slows at a constant rate of 1.0 m/s2 until it stops.


(a) How much time elapses from start to stop (in seconds)?

(b) How far does the vehicle travel from start to stop (in meters)?
Answer #1
Find time taken to reach 22 m/s
Find time taken to reach 0 m/s

10+22=32 s.
The vehicle's first displacement:

The vehicle's second displacement:
Total = 110+242=352m